OpenStudy (anonymous):

A balloon is at a height of 50 meters, and is rising at the constant rate of 5 m/sec. a bicyclist passes beneath it, traveling in a straight line at the constant speed of 10 m/sec. How fast is the distance between the bicyclist and the balloon increasing 2 seconds later?

OpenStudy (anonymous):

What is the formula for the distance between the bicyclist and the balloon?

OpenStudy (anonymous):

Oooh, fun one =)

OpenStudy (anonymous):

I find it always helps to draw a picture..

OpenStudy (anonymous):

Sqrt(10^2+5^2)

OpenStudy (anonymous):

I drew a right triangle with the balloon on top and the bike on the right acute corner

OpenStudy (anonymous):

think of it as if the balloon were stationary

OpenStudy (anonymous):

Refer to picture attached

OpenStudy (anonymous):

:D looks like mine.

OpenStudy (anonymous):

11.180339887498948482045868343656

OpenStudy (anonymous):

Write an equation for d.

OpenStudy (anonymous):

h=sqrt(50^2+y^2)

OpenStudy (anonymous):

=) I'm sure you're more artistic than me though.

OpenStudy (anonymous):

d sorry, I labeled it h

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

y is one leg of the triangle, and it starts at 50, but we're not going to worry about that yet. The other leg is x.

OpenStudy (anonymous):

oh boo.

OpenStudy (anonymous):

after time t seconds, the cyclist will have travelled 2t meters horizontally away from the balloon. the balloon would have risen 50 + 10t meters vertically away from the cyclist. the distance between them is sqrt((2t)^2 + (50 + 10t)^2) = sqrt(104t^2 + 2500 + 1000t) rate of change of distance = (208*t + 1000)/2sqrt(104t^2+2500+1000t) = (104t + 500)/sqrt(104t^2 + 2500 + 1000t) at t = 4, rate of change = (104*4 + 500)/sqrt(104*16 + 2500 + 4000)

OpenStudy (anonymous):

so d=sqrt(x^2+y^2)

OpenStudy (anonymous):

Yes. Exactly.

OpenStudy (anonymous):

ok then differentiate that?

OpenStudy (anonymous):

for x?

OpenStudy (anonymous):

Well, it's changing two ways. Let's start with dd/dx

OpenStudy (anonymous):

hmm ok.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

dh/dx=(x+y dy/dx)/sqrt(x^2+y^2)

OpenStudy (anonymous):

Oh, okay. I think I have a simpler way.

OpenStudy (anonymous):

Mind if we change our approach?

OpenStudy (anonymous):

no worries.

OpenStudy (anonymous):

First, write x as a function of t, time. Also write y as a function of t, time.

OpenStudy (anonymous):

dx/dt=5 and dy/dt=10

OpenStudy (anonymous):

I was hpoing there was a simpler way since we had that info

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Isn't it true that x = 10t?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

And y?

OpenStudy (anonymous):

oh eek. no sure.

OpenStudy (anonymous):

What does y equal in terms of t?

OpenStudy (anonymous):

It starts at 50.

OpenStudy (anonymous):

And increases at a rate of 5m/s.

OpenStudy (anonymous):

wait, I'm confused. x is height right?

OpenStudy (anonymous):

how does it start at 50? because it passes under the balloon?

OpenStudy (anonymous):

No, x is the horizontal distance, the distance the bike has traveled.

OpenStudy (anonymous):

y is the vertical distance, the distance of the balloon from the ground, and it starts at 50 m high.

OpenStudy (anonymous):

y=5t

OpenStudy (anonymous):

+50

OpenStudy (anonymous):

Yes. Exactly. y = 50 + 5t.

OpenStudy (anonymous):

Okay, so x = 10t. y = 50 + 5t. Also, we got that d = sqrt(x^2 + y^2)

OpenStudy (anonymous):

Follow?

OpenStudy (anonymous):

so we just plug it in?

OpenStudy (anonymous):

oh goddamn, i misread y = 50 + 5t, x = 2t distance = sqrt(2500 + 500t + 29t^2) d(distance)/t = (500 + 58t)/2sqrt(2500 + 500t + 29t^2) = (250 + 29t)/sqrt(2500 + 500t + 29t^2) at t = 2: d(distance)/t = (250 + 58)/sqrt(2500 + 1000 + 116)

OpenStudy (anonymous):

Yes. We substitute for x and y.

OpenStudy (anonymous):

So d = sqrt((10t)^2 + (50 + 5t)^2)

OpenStudy (anonymous):

ok. then do it like an mx+b problem? eek

OpenStudy (anonymous):

No not quite.

OpenStudy (anonymous):

It's going to be a derivative problem. But first simplify.

OpenStudy (anonymous):

oh ok wonderful

OpenStudy (anonymous):

I'll let you simplify that.

OpenStudy (anonymous):

ok h=sqrt(125t^2+500t+2500)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I'll simplify it some more. give me a sec

OpenStudy (anonymous):

That's as far as you need to simplify it.

OpenStudy (anonymous):

So what is the question asking for? dh/dt, right?

OpenStudy (anonymous):

oh ok. so just take the derivative

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

k so dh/dt=(125t+250)/sqrt(125t^2+500t+2500)

OpenStudy (anonymous):

Yes, I believe.

OpenStudy (anonymous):

You're beautiful, btw.

OpenStudy (anonymous):

oh thanks. how do you know which one I am? :P

OpenStudy (anonymous):

I don't. They're both beautiful.

OpenStudy (anonymous):

hahaha. good answer.

OpenStudy (anonymous):

ok, so after I take the derivative, I still have t in htere as a variable. :(

OpenStudy (anonymous):

What does the question ask?

OpenStudy (anonymous):

oh psh, I'm a dunce. sorry

OpenStudy (radar):

Post the answer, I am trying to work this also.

OpenStudy (anonymous):

Haha don't be sorry.

OpenStudy (anonymous):

it asks for the value of the derivative when t = 2

OpenStudy (anonymous):

ok got it. 500/sqrt(4000)

OpenStudy (radar):

About 8 meters/sec

OpenStudy (anonymous):

yes.

OpenStudy (radar):

Do you have the approved answer?

OpenStudy (anonymous):

no, I don't

OpenStudy (radar):

smoothmath for our edification is that a correct answer

OpenStudy (anonymous):

Good work, Nasryn. =)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

It is the right answer.

OpenStudy (radar):

I take it a medal is earned, thanks for the step by step and lucid explanation and drawing.

OpenStudy (anonymous):

finally on my last one! wonderful! I might get some sleep tonight. :P

OpenStudy (anonymous):

Haha. I appreciate it, Radar. Math is it's own reward though. =)

OpenStudy (radar):

You are so right !!

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

Yay, Nasryn! Don't forget to have fun too.

OpenStudy (anonymous):

jesus christ, if she's the one on the right, she's smoking.

OpenStudy (anonymous):

if she's the one on the left, still hot.

OpenStudy (anonymous):

I'm the one on the right.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

yeah...don't mind the math geeks as we drool for a while our saliva descends at 2 meters/sec downwards

OpenStudy (anonymous):

Yeah. Freaking beautiful. Mind if I facebook you?